Assume that the traffic to the web site of Smiley’s People, Inc., which sells customized T-shirts, follows a normal distribution, with a mean of 4.46 million visitors per day and a standard deviation of 800,000 visitors per day.

(b) What is the probability that the web site has 3 million or more visitors in a single day? If needed, round your answer to four decimal digits.
Answer #1

Normal distribution: P(X < A) = P(Z < (A - mean)/standard deviation)

Mean = 4.46 million visitors per day

Standard deviation = 0.8 million visitors per day

b) P(the web site has 3 million or more visitors in a single day), P(X > 3)

= 1 - P(X < 3)

= 1 - P(Z < (3 - 4.46)/0.8)

= 1 - P(Z < -1.825)

= 1 - 0.0340

= 0.9660